INAF LBC+MODS queue observing, 2013 January 11-12

Observer: M.Faccini, R. Speziali, F.Cusano
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: M. Edwars
Instrument: LBC+MODS

Summary

Just few clouds at west horizon at sunset, sky is clear. Dewpoint at -16\xB0C and external temperature at - 15\xB0C
There is a difference about 12/13 degres between primary mirrors and external temperature
We have to wait changes in weather conditions.
After few hours the external temperature falls down to -19\xB0C, the mirrors temperature is still far from the external temperature. At 3:30 the TO decides to not open the dome for this night ( we are almost at -20\xB0C)

Data Summary

Project Objects Filter ExpTime Notes
         

Details

-- MarcoFaccini - 12 Jan 2013
Topic revision: r3 - 12 Jan 2013, MarcoFaccini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback